Home‎ > ‎Crew 133‎ > ‎

112-ScienceReport

Science Report
01/12/2014
Paula Crock
 
Today was the second day of data collection for the EVA helmet temperature study and the radio 'dead'
zone mapping survey.
 
EVA helmet temps:  Data were again collected on two subjects during today's EVA.  This time the
helmet sensors were taped at the back of the helmet to keep them from coming out, which solved the
problem experienced yesterday.  Otherwise the placement method seems to work well.
 
Radio Dead Zone map:  For this EVA today we headed south east and then east to get data on two of
Jessup's strata study.  During those excursion I completed 5 radio comms checks when we were in
particularly topologically complex regions.  For future EVA, unless the plan is to check specific
model-derived map boundaries comms checks would best be performed either at long line of sight
distances from the Hab or in topologically complex regions as they are encountered rather than
trying to determine them in advance from maps.

-----------------

Science Report
01/12/2014
Joseph Jessup


Geology Report

 Acquired two points for GIS map study of canyon drainage flow patterns via EVA on foot.  Measured
strata elevations for ARC GIS software using GPS and physical ground truthing and climbing up hill
slopes  at:
 12S 518343 4250 580 Elevation 4493  and 12S 518 676 4250 860 Elevation 4488

 Two more elevation points are needed, and data collection for mapping project will be complete.

 ARC GIS 10.0 strata reconstruction map 50% completed.